1. Apply your mathematical models to your allocated car. Use the given data for the 0 – 28 m/s and 400m times to calculate the:




1. value of the coefficient A



2. maximum velocity



3. maximum acceleration.




Given: t (0-28 m/s) is 1.9s, t (400m) is 10.50s & tmaxspeed is 7.1s.



Given velocity equation :V (t) = A (1 - e^((-t)/(t maxspeed)))




i need step by step solution to understand




1
Expert's answer
2022-05-30T16:06:10-0400

 

"v(t) = A (1-e ^ {- t\/t_{maxspeed}})"


Using the information that "t_{maxspeed}=7.1\ s" we have


"v(t) = A (1-e ^ {- t\/7.1})"

Using the information that "t (0-28 m\/s)" is "1.9\ s" we have


"v(1.9) = A (1-e ^ {- 1.9\/7.1})=28"

"A=\dfrac{28}{1-e ^ {- 1.9/7.1}}m/s"

"A=119.255\ m/s"

2.


"v_{max}=v(t_{maxspeed})=A (1-e ^ {- t_{maxspeed}\/t_{maxspeed}})"

"v_{max}=A(1-e^{-1})"

"v_{max}=119.255\ m/s\cdot(1-e^{-1})"

"v_{max}=75.387\ m/s"

3.


"a(t)=\dfrac{dv}{dt}=\dfrac{A}{t_{maxspeed}} (e ^ {- t/t_{maxspeed}})"

"=\dfrac{119.255\ m/s}{7.1\ s} (e ^ {- t/7.1}), 0\ s\le t\le7.1\ s"

The function "a(t)" decreases for "0\ s\le t\le7.1\ s."


"a_{max}=a(0)=\dfrac{119.255\ m/s}{7.1\ s}"

"a_{max}=16.797\ m/s^2"